# “Rich-Get-Richer”? Analyzing Content Creator Earnings Across Large Social Media Platforms

### v1

@@ 2.3  Power Laws and Algorithmically-Driven ‘Rich-Get-Richer’ Dynamics
where $P(A)$ is the probability of observing a node (e.g., a content creator) with attention $A$ (such as clicks, views, or follows) greater than or equal to a certain threshold (the cumulative distribution function, CDF), and the exponent $\alpha$ characterizes the degree of inequality in how attention is distributed across the network.
Power law distributions appear in many domains [28], often driven by underlying self-reinforcing processes. One such process is captured by the model introduced by [6], which effectively illustrates how compounding algorithmic advantage can emerge in social and digital networks. In online platforms, a small number of creators can accumulate disproportionate attention over time, aided by algorithms that recommend popular or trending content more aggressively.
The foundational mechanism is akin to a reinforced Pólya urn model, adapted here to the digital attention landscape. As a network of creators and consumers grows, new attention events – new views, clicks, or follows – do not occur uniformly at random. Instead, they tend to cluster around creators who already have substantial visibility. This reinforces the popularity of a few, resulting in a ‘rich-get-richer’ dynamic commonly observed on platforms governed by algorithmic curation and recommendation [5, 41, 28, 63, 21, 7, 8].
To model this, we introduce a parameter $\gamma$ that balances between random attention allocation (exploration) and attention guided by existing popularity (exploitation):
  • With probability $\gamma$, an attention event (e.g., a new user’s click, view, or follow) is directed to a creator chosen uniformly at random, independent of current popularity.
  • With probability $1-\gamma$, the likelihood of directing attention to a particular creator is proportional to that creator’s existing level of attention.
This parameter $\gamma$ represents how the platform’s algorithm navigates between exploration (showing less-known creators, broadening the range of content surfaced) and exploitation (showing widely-popular creators, reinforcing established winners). A low $\gamma$ implies strong preferential attachment to already-popular creators, amplifying their advantage, whereas a high $\gamma$ distributes attention more evenly and blunts the rich-get-richer effect.
Originally, the Barabasi-Albert (BA) model assumed full preferential attachment (equivalent to $\gamma=0$) [6], leading directly to a power law exponent $\alpha=2$.55endnote: 5See Appendix B.2 for the original derivation. Here, we generalize that approach to show how $\alpha$ depends on the exploration parameter $\gamma$.
By extending the BA model and following a heuristic derivation,66endnote: 6See Appendix B for details. we find:
@@
In this formulation, reducing $\gamma$ (i.e., relying more on popularity-driven exploitation) decreases $\alpha$, intensifying inequality and concentrating attention among top creators. Conversely, increasing $\gamma$ promotes exploration, raising $\alpha$ and yielding a more balanced distribution of attention.
@@ 2.4  Interpretation of $\alpha$
The exponent $\alpha$ provides insights into the structure of algorithmically mediated attention ecosystems. Since $\alpha$ is directly shaped by $\gamma$, it reflects how the platform’s algorithmic preferences influence the overall distribution of attention:
  • A smaller $\alpha$ (closer to 2) indicates that the algorithm heavily favors already-popular creators (lower $\gamma$), leading to a steeply skewed distribution of attention.
  • A larger $\alpha$ corresponds to a more equitable distribution of attention, arising from a higher $\gamma$, i.e., an algorithm that is more “exploratory” and less biased toward existing popularity.
This approach mirrors the logic of a Pólya urn with reinforcement [14, 15], where the probability of adding a ball of a certain color increases with the number already in the urn. In the digital realm, each new click or view (analogous to adding a ball) is more likely to go to a popular creator (a “color” that already has many balls). Adjusting $\gamma$ changes the algorithm’s propensity to reinforce existing popularity.
Applied to platform earnings, the same attention-reinforcing mechanisms shape creators’ revenue streams. If visibility and engagement are governed by these dynamics, then earnings—derived from ads, sponsorships, or subscriptions—also become concentrated among a few popular creators. By estimating $\alpha$ from observed earnings distributions, we can quantify how deeply entrenched the rich-get-richer effect is in these digital ecosystems and assess whether algorithmic changes (modifying $\gamma$) can create more sustainable conditions for a broader “middle class” of creators.
@@ 3  Results
To fit the power law tail, we apply the method introduced by [16], which relies on maximum likelihood estimation to fit the power law model and employs the Kolmogorov-Smirnov (KS) statistic to assess the goodness-of-fit. This approach estimates the scaling parameter by maximizing the likelihood of the observed data and determines the lower bound for power law behavior by minimizing the distance between the empirical and theoretical distributions. The analysis is conducted using the poweRlaw package [32] in R [54].
@@
Figure 1 displays the log-log complementary cumulative distribution function (CCDF) of earnings data across the major social media platforms YouTube, Instagram, Twitch, Patreon, Twitter, and Facebook, along with the ideal power law lines for each in red (sloping downwards). The red vertical dotted line represents the estimated minimum value above which the earnings distribution adheres to a power law pattern. Several key findings emerge.77endnote: 7See Appendix A for a discussion on the robustness of the results.
@@
The earnings distributions for most platforms in Figure 1 exhibit a strong power law pattern around $\alpha=2$.88endnote: 8A power law exponent of 2 means that as you move to higher earnings, the probability of observing such high earners decreases as the square of their earnings, formally $P(X>x)\sim\frac{1}{x^{\alpha}}=\frac{1}{x^{2}}$. For example, if 1,000 creators earn \$100, only about 100 will earn \$1,000, and only 10 will earn \$1,000,000. This follows from the scale invariance properties of the power law coupled with our finding that the power law begins very early on in the earnings distribution. This means that the benefits to moving up the algorithmic ‘ladder’ are likely enormous at each rung, highlighting just how important it is to be higher up the screen, or the user’s attention sphere. Position matters enormously to monetization. YouTube, Instagram, Twitch, and Facebook consistently show well-defined power law distributions, ranging between 1.8 and 1,94 for the year 2023, suggesting significant earnings concentration among top creators. Patreon and Twitter show a much weaker power law (2.24 and 2.35, respectively).
@@
Power law dynamics emerge early at between \$26 and \$70 per monthly earnings across these platforms, a very low monthly earnings amount consistent across platforms. At an annual amount of \$312 to \$840 this is in stark contrast to standard income distributions, where the power law typically applies to income levels only above \$100,000 per annum [20, 61].
@@
Unlike income distribution models where the power law typically governs only the extreme right tail [61],99endnote: 9With the majority of the distribution better explained by exponential or gamma-type distributions. our results show that social media earnings data are largely dominated by a power law from very early on in the earning distribution. It is worth noting though that the extreme right tail is not as concentrated as a pure power law for most platforms, represented by the grey data points in Figure 1 falling off below the slanted red power law lines. Twitter, however, shows a strong power law on its right-most tail, implying that earnings become extremely concentrated but only once a higher threshold is reached (\$2,114). This means that opportunities for Patreon content creators on Twitter are likely to be greater than other comparable platforms below this threshold (especially for ‘NSFW’ creators). Figure 2 plots the average $\alpha$ by platform (left hand side) and year (right hand side).
@@
Platform-specific power law exponents ($\alpha$) exist and reflect differences in earnings inequality and platform dynamics. YouTube and Twich have the lowest, i.e. most extreme, estimated $\alpha$ values, suggesting relatively heavier tails and higher inequality, indicating a greater tendency for the emergence of super-hubs—nodes with degrees larger than expected in the standard preferential attachment model [6]. Twitter and Patreon consistently display the highest (least extreme) $\alpha$ values, reflecting lighter tails and less earnings inequality.
Furthermore, temporal trends in $\alpha$ show a worsening in the platform power-relation, perhaps as algorithmic power grows. Across Facebook, Instagram, Twitch, and Youtube, $\alpha$ values have generally worsened over time (getting smaller), reflecting a transition toward heavier tails and higher earnings inequality at the extremes, as a pure Pólya urn effect emerges in line with the preferential attachment model [6]. Youtube has the lowest value in 2023, followed closely by Instagram, with values around 1.8 and 1.88. Instagram shows the most extreme change in $\alpha$ values over time, and in particular between 2021 and 2023, which may reflect the large changes made to its platform to promote less social graph and more purely algorithmic content in users’ feeds. Lastly, it is worth noting that Patreon content creators on Twitter appear not to have suffered from Elon Musk’s takeover of the platform in October 2022, with their earnings distribution becoming less unequal (higher $\alpha$) since 2021. Though ‘NSFW’ (not safe for work) creators increased on the platform by almost 30 percentage points, going from 45% and 46% of Patreon creators on Twitter in 2018 and 2021, respectively, to 58% in 2014, reflecting a growing permissiveness on the platform and Instagram cracking down (Table 2 in Appendix).
@@
Figure 3 shows a clear association between the power law exponent ($\alpha$) and the median of the earnings distribution: platforms with higher median earnings tend to exhibit less heavy-tailed distributions (higher $\alpha$). This indicates that high earnings for the middle layer of creators (“middle-class”) are associated with less unequal platforms. Twitter – and to some extent Patreon – stands out in this context, combining a high median income with a relatively equal earnings distribution up until a certain point. This suggests a more robust middle-class of creators, even though Twitter’s dominance of top earners is in fact worse than other platforms. In contrast, platforms with lower $\alpha$ show a “rich-get-richer” pattern far earlier one, where a small elite captures most of the income, leaving the middle-class of creators under-rewarded.
Figure 4 notes the proportion of observations for each platform (pooled across years) which follow a power law distribution. This allows us to assess the extent to which the power law distribution characterizes content creators’ earnings. It shows that the proportion of earnings following a power law distribution varies significantly across platforms, from 5% (Twitter) to almost 55% (Twitch). Twitch and YouTube demonstrate the highest power law proportions (0.55 and  0.47, respectively), suggesting that a large share of creator earnings is dominated by the power law distribution. This may reflect the platforms’ strong algorithmic recommendation systems and high reliance on ‘viral’ or popular content. Instagram and Facebook also show a relatively high and similar proportion (0.40-0.43). Patreon and Twitter show moderate to lower power law proportions, suggesting that only the very top of the income distribution aligns with the power law, as is often observed in standard income distributions [20, 61].
@@ 4.1  Restatement of core findings
By detailing the power law dynamics governing digital creator economies, this paper contributes to broader discussions on fairness, sustainability, and innovation in networks [9]. Our findings highlight the central role of algorithmic design in shaping earnings inequality and emphasize the need for interventions that work to moderate algorithmic effects, balancing fairness with platform engagement objectives. A key limitation is the use of Patreon content creators as a proxy for the health of platform content creators themselves, given that the underlying monetization mechanisms do differ somewhat.

### v2

@@ 2.3  Power Laws and Algorithmically-Driven ‘Rich-Get-Richer’ Dynamics
where $P(A)$ is the probability of observing a node (e.g., a content creator) with attention $A$ (such as clicks, views, or follows), and the exponent $\alpha$ characterizes the degree of inequality in how attention is distributed across the network.
Power law distributions appear in many domains [gabaix2009power], often driven by self-reinforcing processes. One such process is captured by the model of [barabasi1999emergence], which illustrates how compounding algorithmic advantage emerges in social and digital networks. The foundational mechanism resembles a reinforced Pólya urn model [chung2003generalizations, chung2006concentration], where the probability of adding a ball of a certain color increases with the number already in the urn. In the digital realm, each new click or view is more likely to go to creators who already have substantial visibility, resulting in a ‘rich-get-richer’ dynamic commonly observed on platforms governed by algorithmic curation [barabasi2005origin, jackson2008social, gabaix2009power, vidal2011interactome, easley2010networks, barabasi2011network, barbosa2018human].
To model this, we introduce a parameter $\gamma$ that governs the balance between random algorithmic attention allocation (exploration) and attention guided by existing popularity (exploitation):
  • With probability $\gamma$, an attention event (e.g., a new user’s click, view, or follow) is directed to a creator chosen uniformly at random.
  • With probability $1-\gamma$, attention is directed proportionally to a creator’s existing attention level.
This parameter captures how platform algorithms trade off exploration (surfacing less-known creators) against exploitation (reinforcing established winners). A low $\gamma$ implies strong preferential attachment to already-popular creators, whereas a high $\gamma$ distributes attention more evenly and weakens the rich-get-richer effect.
In the original Barabási–Albert (BA) model there is no random component: new links attach purely in proportion to existing degree, which generates a power-law degree distribution with density exponent $\alpha=3$.1010endnote: 10In this mixed model, choosing $\gamma=1/2$ yields the same exponent $\alpha=3$, although the underlying mechanism differs from the original BA construction. See Appendix B.2 for the standard BA derivation. A closely related BA-style mixed (copying) model, in which attention is allocated according to the two rules above, yields a power-law density $p(A)\propto A^{-\alpha}$ with1111endnote: 11See Appendix B for more details.
@@
Within this mixed model, reducing $\gamma$ (relying more on popularity-driven exploitation) lowers $\alpha$, intensifying inequality and concentrating attention among top creators, while increasing $\gamma$ raises $\alpha$ and produces a more balanced distribution of attention.
@@ 2.4  Interpretation of $\alpha$
The exponent $\alpha$ provides insights into algorithmically mediated attention ecosystems. Since $\alpha$ is directly shaped by $\gamma$, it reflects how platform algorithmic preferences influence overall attention distribution:
  • A smaller $\alpha$ (closer to 2) indicates algorithms heavily favor already popular creators (lower $\gamma$), leading to steeply skewed attention distributions.
  • A larger $\alpha$ corresponds to more equitable attention distribution, arising from higher $\gamma$ algorithms that are more exploratory and less biased toward existing popularity.
Applied to platform earnings, the same attention-reinforcing mechanisms shape creators’ revenue streams. If visibility and engagement are governed by these dynamics, then earnings – including earnings derived from ads, sponsorships, or subscriptions – also become concentrated among a few popular creators. By estimating $\alpha$ from observed earnings distributions, we can quantify how deeply entrenched the rich-get-richer effect is in these digital ecosystems and assess whether algorithmic changes (modifying $\gamma$) can create more sustainable conditions for a broader “middle class” of creators.
@@ 3  Results
To fit the power law tail, we apply the method introduced by [clauset2009power], which relies on maximum likelihood estimation to fit the power law model and employs the Kolmogorov-Smirnov (KS) statistic to assess the goodness-of-fit. This approach estimates the scaling parameter by maximizing the likelihood of the observed data and determines the lower bound for power law behavior by minimizing the distance between the empirical and theoretical distributions. The analysis is conducted using the poweRlaw package [poweRlaw] in R [R_core_team_2024].
@@
Figure 1 displays the log-log complementary cumulative distribution function (CCDF) of earnings data across the major social media platforms YouTube, Instagram, Twitch, Patreon, Twitter, and Facebook, along with the ideal power law lines for each in red (sloping downwards). The red vertical dotted line represents the estimated minimum value above which the earnings distribution adheres to a power law pattern. Several key findings emerge.1212endnote: 12See Appendix A for a discussion on the robustness of the results.
@@
The earnings distributions for most platforms in Figure 1 exhibit a strong power law pattern around $\alpha=2$.1313endnote: 13Formally, if individual earnings $X$ follow a Pareto-type distribution with density $p(x)\propto x^{-\alpha}$ for $x\geq x_{\min}$ and $\alpha>1$, then the upper tail behaves as $P(X\geq x)\propto x^{-(\alpha-1)}$. Thus, for $\alpha\approx 2$, the probability of observing earnings above a level $x$ falls roughly inversely with $x$: increasing earnings by a factor of 10 makes such high earners about 10 times rarer. For example, if 10,000 creators earn at least \$100, we would expect around 1,000 to earn at least \$1,000, about 100 to earn at least \$10,000, and roughly 1 to earn at least \$1,000,000. This follows from the scale invariance properties of the power law coupled with our finding that the power law begins very early on in the earnings distribution. This means that the benefits to moving up the algorithmic ‘ladder’ are likely enormous at each rung, highlighting just how important it is to be higher up the screen, or the user’s attention sphere. Position matters enormously to monetization. YouTube, Instagram, Twitch, and Facebook consistently show well-defined power law distributions, ranging between 1.8 and 1.94 for the year 2024, suggesting significant earnings concentration among top creators. Patreon and Twitter show a much weaker power law (2.24 and 2.35, respectively).
@@
Power law dynamics emerge early at between \$26 and \$70 per monthly earnings across these platforms, a very low monthly earnings amount consistent across platforms. At an annual amount of \$312 to \$840 this is in stark contrast to standard income distributions, where the power law typically applies to income levels only above \$100,000 per annum [druagulescu2001exponential, tao2019exponential].
@@
Unlike income distribution models where the power law typically governs only the extreme right tail [tao2019exponential],1414endnote: 14With the majority of the distribution better explained by exponential or gamma-type distributions. our results show that social media earnings data are largely dominated by a power law from very early on in the earning distribution. It is worth noting though that the extreme right tail is not as concentrated as a pure power law for most platforms, represented by the grey data points in Figure 1 falling off below the slanted red power law lines. Twitter, however, shows a strong power law on its right-most tail, implying that earnings become extremely concentrated but only once a higher threshold is reached (\$2,114). This means that opportunities for Patreon content creators on Twitter are likely to be greater than other comparable platforms below this threshold (especially for ‘NSFW’ creators). Figure 2 plots the average $\alpha$ by platform (left hand side) and year (right hand side).
@@
Platform-specific power law exponents ($\alpha$) exist and reflect differences in earnings inequality and platform dynamics. YouTube and Twitch have the lowest, i.e. most extreme, estimated $\alpha$ values, suggesting relatively heavier tails and higher inequality, indicating a greater tendency for the emergence of super-hubs—nodes with degrees larger than expected in the standard preferential attachment model [barabasi1999emergence]. Twitter and Patreon consistently display the highest (least extreme) $\alpha$ values, reflecting lighter tails and less earnings inequality.
Furthermore, temporal trends in $\alpha$ show a worsening in the platform power-law relation, perhaps as algorithmic power grows. Across Facebook, Instagram, Twitch, and Youtube, $\alpha$ values have generally worsened over time (getting smaller), reflecting a transition toward heavier tails and higher earnings inequality at the extremes, as a pure Pólya urn effect emerges in line with the preferential attachment model [barabasi1999emergence]. Youtube has the lowest value in 2024, followed closely by Instagram, with values around 1.8 and 1.88. Lastly, it is worth noting that Patreon content creators on Twitter appear not to have suffered from Elon Musk’s takeover of the platform in October 2022, with their earnings distribution becoming less unequal (higher $\alpha$) since 2021. Though ‘NSFW’ (not safe for work) creators increased on the platform by almost 30 percentage points, going from 45% and 46% of Patreon creators on Twitter in 2018 and 2021, respectively, to 58% in 2024, reflecting a growing permissiveness on the platform (Table 2 in Appendix).1515endnote: 15Note that these time-series patterns should be interpreted cautiously: the sample consists of only three cross-sections (2018, 2021, 2024), and a few platform-year estimates, especially Patreon and Instagram in the final year, are somewhat sensitive to the estimation method used (see Appendix A).
@@
Figure 3 shows a clear association between the power law exponent ($\alpha$) and the median of the earnings distribution: platforms with higher median earnings tend to exhibit less heavy-tailed distributions (higher $\alpha$). This indicates that high earnings for the middle layer of creators (“middle-class”) are associated with less unequal platforms. Twitter – and to some extent Patreon – stands out in this context, combining a high median income with a relatively equal earnings distribution up until a certain point. This suggests that Twitter has a more robust middle class of creators. In contrast, platforms with lower $\alpha$ show a “rich-get-richer” pattern far earlier one, where a small elite captures most of the income, leaving the middle-class of creators under-rewarded.
Figure 4 notes the proportion of observations for each platform (pooled across years) which follow a power law distribution. This allows us to assess the extent to which the power law distribution characterizes content creators’ earnings. It shows that the proportion of earnings following a power law distribution varies significantly across platforms, from 5% (Twitter) to almost 55% (Twitch). Twitch and YouTube demonstrate the highest power law proportions (0.55 and  0.47, respectively), suggesting that a large share of creator earnings is dominated by the power law distribution. This may reflect the platforms’ strong algorithmic recommendation systems and high reliance on ‘viral’ or popular content. Instagram and Facebook also show a relatively high and similar proportion (0.40-0.43). Patreon and Twitter show moderate to lower power law proportions, suggesting that only the very top of the income distribution aligns with the power law, as is often observed in standard income distributions [druagulescu2001exponential, tao2019exponential].
@@ 4  Discussion: Market Structure, Algorithms, & Policies
By detailing the power law dynamics governing digital creator economies, this paper contributes to broader discussions on fairness, sustainability, and innovation in networks [bianconi2001competition]. Our findings highlight the central role of algorithmic design in shaping earnings inequality and emphasize the need for interventions that work to moderate algorithmic effects, balancing fairness with platform engagement objectives.
4.1  Platform market power and inequality in attention & earnings
The relationship between platforms’ algorithmic choices over what content to prioritize and their market power has been theorized in the context of advertising and data privacy [srinivasan2019antitrust, strauss2024Amazon, o2024algorithmic].1616endnote: 16Market power reflects how competitive the social media market is—also known as the underlying ‘market structure’—defined by the number of firms in the market and shaped by cost structure and technology [carlton2015modern].
Similar dynamics appear to shape content creator earnings. In our data, Patreon-only and Twitter-only creators have more equal earnings distributions than those who monetize via more dominant platforms such as YouTube or Instagram. One interpretation is that less concentrated market structures are associated with less skewed algorithmic allocation of attention and earnings.
However, greater competition in the social media market has had ambiguous effects on the health of the ecosystem, including for content creators. Competition from TikTok appears to have locked the market into highly addictive, engagement-maximizing algorithms – a race to the bottom. TikTok accelerated the shift away from the social graph towards purely algorithmic selection of content. While TikTok may promote more long-tail content, we lack evidence on whether this has translated into a more equal distribution of earnings within the platform, given the competitive pressures it faces [brennan2020attention].
At the same time, competition has raised total payouts to creators as platforms bid for their participation [FT2023YouTubeShorts]. Yet this may increase within-platform inequality if top-tier creators are the main beneficiaries of competitive poaching. It also highlights that creators do not always multi-home to the extent often assumed. Many focus on a single platform, as in our sample, so competition between platforms is only imperfectly transmitted to them, even if consumers do multi-home. In this context, greater interoperability – via shared, open protocol infrastructures that allow creators to take followers with them across platforms – may do more to reshape earnings dynamics than competition from platform entry alone.
So long as platforms retain strongly power-law–promoting algorithms, and monetize on that same basis, earnings inequality is likely to persist. The more followers and views a creator has, the more they can earn [Naceva_2025_CreatorEP]. Gains to stardom are multiplicative rather than linear: outsized brand deals and advertising revenue shares accrue to a small group of superstar creators with enormous follower counts. It is notable that YouTube offers one of the most generous shares of ad sales, and hosts the greatest earnings inequality among Patreon creators who monetize there [FT2023YouTubeAdSharing].
More competition alone is, therefore, unlikely to correct these unequalizing algorithms, especially if they dampen rational consumer choice and the usual benefits of competition on the user side. Interoperability mandates on the developer and creator side are likely a more promising route to raising the floor in creator earnings then, by making it easier for creators to move between competing platforms without leaving their audience behind.
A central challenge facing recommender systems on mature platforms is their potential to compound initial advantage (viewership), by promoting popular, existing material and focusing less on untested ‘fresh’ content [covington2016deep, sutton2018reinforcement]. To the extent that this becomes baked into a platform’s business model, with associated commercial incentives, it can become difficult to dislodge through market forces alone.
A “longer-tail” algorithmic system [anderson2012long], whereby a more diverse range of content is explored and promoted, invariably trades off some (more) immediate profit (exploitation of existing content), in favor of greater content exploration [hu2023incentivizing, yao2024unveiling]. But maximizing for immediate user engagement [immorlica2024clickbait], as a quarterly economic imperative facing the platform’s shareholders, can lead to less new content being created and even less longer-term user engagement – even if potentially at the expense of diminished immediate user satisfaction [facebook_notifications_2021, yao2024unveiling]. Several potential reforms are possible to grow healthier – more diverse and equitable – ecosystems of content promotion and consumption:
Firstly, it remains to be seen if a new entrant into the market could change what the market is currently optimizing for – for the better. Although the competitive dynamics of the market have seen a race to abandon the social graph and maximize time spent on platform through short-form videos, users also show a preference for longer-form content too as evidenced by the rise of lengthy podcasts. If the market needs a push in this direction then regulations that limit infinite scroll mechanisms would be one means of incentivizing algorithmic optimizations for longer-form content. But it is still unclear if this would benefit content creators and make for algorithmic recommendations that are more equalizing – that tap into the long-tail more effectively.
Secondly, we recommend internal company measures that incentivize long-term orientated (algorithmic) A/B testing, to help highlight the business case for longer-term business optimizations at the algorithmic level [mazzucato_strauss_2024]. [hu2023incentivizing] recommends modifying the learning rates of recommendation algorithms to help balance the trade-off between exploration and exploitation, ensuring that high-quality content is recognized and then promoted quickly. Recent research suggests that incorporating mechanisms to surface underrepresented (“longer-tail”) content [anderson2012long] could enhance overall platform performance without necessarily sacrificing user satisfaction [houtti2024leveraging, yao2024unveiling].
Use of longer-tail discovery algorithms on social media to unlock and match a greater variety of non-viral content to users might also be incentivized through mandatory platform mechanisms that penalize the creation of low-quality content from dominant creators [hu2023incentivizing]. Facebook has recently chosen to do the opposite [techcrunch_2025].
Thirdly, greater platform disclosure of the drivers of platforms’ recommendation systems, in-line with the Digital Services Act (DSA, Article 27) of the European Union – may help with the adoptions of long-tail recommender systems [park2008long]. Platforms should be required to disclose key details about their recommendation algorithms to regulators, researchers, and the public. These disclosures should include the algorithmic weighting of past engagement metrics, such as views and likes, which often amplify the visibility of previously successful content. In addition, platforms should clarify the criteria by which new or less popular content is surfaced within the recommendation engine, ensuring greater accountability for algorithmic bias. In addition, mechanisms in place that punish low quality, click-bait content, should ideally be detailed. In the DSA mandated platform audits, the extent to which AI and ML recommendation systems favor certain types of creators, content genres, or user demographics should be evaluated.
Platforms could allow users to customize their feeds, such as through “content diversity sliders” to adjust the balance between popular and niche content. YouTube, for example, sometimes tries to encourage this by asking if you want to explore unrelated topics to what you have been watching. And although Article 27 of the DSA requires platforms to offer users such “non-profiling” alternatives (e.g. chronological feeds), this is seen as an alternative to algorithmically curated ones.
5  Conclusion